OpenStudy (anonymous):

PLEAS HELP STUCK ON LAST QUESTION OF MY HW find the slope intercept form for the line satisfying the following conditions x-intercept 3, y-intercept 1/3

OpenStudy (tommynaut):

By being given intercepts, you're actually being given the two points (3,0) for the x-intercept and (0, 1/3) for the y-intercept. You could use the gradient formula to find the gradient of the line and that should lead you in the right direction, assuming you know the gradient formula and the slope/intercept form of a line.

OpenStudy (anonymous):

to tell you the truth right now im to tired that i just cant concentrate been doing hw all day if you could help me more it would be great!

OpenStudy (tommynaut):

\[\frac{ y _{2}^{} - y _{1}}{x _{2} - x _{1} }\] That is the gradient formula. If you substitute in the values, that should get you the gradient, which is represented by m in the slope/intercept formula y = mx + b

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

it would be 1/3 divided by 3?

OpenStudy (anonymous):

is it y=-1/9x+1/3

OpenStudy (tommynaut):

You could check for yourself if you substitute in x=0 to discover the y-intercept and y=0 to find the x-intercept. If the intercepts you get match the ones given in the question, then yes, you're right!

OpenStudy (tommynaut):

But yeah, you're right :P Good job

OpenStudy (anonymous):

thanks for the help!!

OpenStudy (tommynaut):

No problem :)